Y=x+2 Find the slope, and the Yintercept
Answer by reviewermath(1029)

Q:
Y=x+2 Find the slope, and the Yintercept
Answer:
In y = mx + b, the slope is m and the y-intercept is b.
Therefore, in y = x + 2, the slope is 1 and the y-intercept is 2.